A pycnometer is a glass flask with a tight fitting glass stopper that is used to obtain the volume of a liquid with high accuracy. A pycnometer has a mass of 25.151 grams when empty, and a mass of 42.481 grams when completely filled with water. The pycnometer has a mass of 47.629 when completely filled with an unknown liquid. What is the density in g/cm³ of the unknown liquid? Assume the density of water is 1.000 g/cm³.
1
Expert's answer
2020-09-17T07:14:03-0400

42.481 g - 25.151 g = 17.330 g - mass of water

it's volume is x cm3

d = m/V

1.000 g/cm³ = 17.330 g / x cm3

x = 17.330 cm3


47.629 g - 25.151 g = 22.478 g - mass of the unknown sample

the volume is 17.330 cm3

d = 22.478 g / 17.330 cm3 = 1.297 g/cm3